OpenStudy (anonymous):

last one! The spinner shown is to be spun 40 times. Which is the best prediction for the number of times the spinner will stop on A or B?

OpenStudy (anonymous):

pic?

OpenStudy (anonymous):

5. The spinner shown is to be spun 40 times. Which is the best prediction for the number of times the spinner will stop on A or B? A. 10 B. 20 C. 25 D. 30

OpenStudy (anonymous):

C

OpenStudy (anonymous):

There are a total of 8 possibilities and a total of 5 "correct" options (2 A + 3B) so we can do \[P(A or B) = \frac{correct}{total}= \frac{5}8{}\] Then the total is \[\frac{5}{8} \times 40 = 25\] C

OpenStudy (anonymous):

thank you guys :D

OpenStudy (anonymous):

Youre Welcome!

OpenStudy (anonymous):

you both gave the right answer who do i give the medel too?